II. 1. Find the eigenvalues and the eigenfunctions for the following Sturm-Liouville problem X+AX=0, x(0)=0, X(TT) = 0
Answer #1

1. X +1 X=0, X(0)=0, x(a)=0 het x=emx (40) be the trial solution of the geven differential equation. The auxilary equation- using X(0)=0 and x(x)=0 we get, A=0 and Bq casqa=0 = 60897 = 0 [i B +0,9+07 997=(2n-1)2, =1,2,.--* 39 = (2n-1), n=1,2,...